Let A be an m x r matrix, B be an m x s matrix, C be an r x n matrix, and D be an s x n matrix. Prove that
(A B) (C
D) -> C on top of D
= AC + BD
1
Expert's answer
2020-02-13T08:56:39-0500

I think there is a mistake in the problem, for example, multiplying the m x r matrix "A" by the m x s matrix "B".

Let "A" be 2x3 matrix and "B" be 2x1 matrix, then the product of "A" by "B" won't be defined:


"\begin{pmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{21} & a_{22} & a_{23}
\end{pmatrix}" *"\begin{pmatrix}
 b_{11} \\
 b_{21}
\end{pmatrix}" "=?"


To multiply matrices, the number of columns of one of them must be equal to the number of rows of another one!

For example,


"\begin{pmatrix}
 a_{11} & a_{12} \\
 a_{21} & a_{22} \\
 a_{31} & a_{32}
\end{pmatrix}" * "\begin{pmatrix}
 b_{11} & b_{12} \\
 b_{21} & b_{22}
\end{pmatrix}" ="\begin{pmatrix}
 a_{11}*b_{11}+ a_{12}*b_{21}& a_{11}*b_{12}+a_{12}*b_{22} \\
 a_{21}*b_{11}+ a_{22}*b_{21}& a_{21}*b_{12}+a_{22}*b_{22} \\
 a_{31}*b_{11}+ a_{32}*b_{21}& a_{31}*b_{12}+a_{32}*b_{22}
\end{pmatrix}"


In matrices "A" and "B" we can't do like this!